OpenStudy (anonymous):

Write a Proportion: The Mariners have won 12 games and lost 21. At the rate, how many wins will they have after playing 162 games? Hint: how many total games they played?

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

@thecookiemonster - please help

OpenStudy (anonymous):

12+21=33 162/33=4.909090909... 4.9*12=58.8 58.8 is about 59. Therefore, the Mariners will win 59 games.

OpenStudy (anonymous):

And, therefore, the Mariners suck at that time.

OpenStudy (anonymous):

Anyway, the answer is 59 out of 162, or 59:162.

OpenStudy (anonymous):

or 59/162

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

ya welcome.